OpenStudy (anonymous):

a boat cross a river with velocity 8 km/h. if the resulting velocity of boat is 10 km/h then the velocity of river

OpenStudy (anonymous):

\[10^2 = 8^2+v_{river}^2\] \[100 - 64 = v_{river}^2\] \[6^2 = v_{river}^2\] \[v_{river} = 6\]